OpenStudy (anonymous):

find dy/dx by implicit differentiation: xsiny+ysinx=1

OpenStudy (anonymous):

\[\sin(y)+x\cos(y)y'+y'\sin(x)+y\cos(x)=0\] via the product rule solve for \(y'\)